# Systematical study of exotic $1^{-+}$ tetraquark spectroscopy

### v1

@@ Abstract
The masses of exotic quantum-number $1^{-+}$ compact tetraquark states are calculated in a constituent quark model (CQM), where Cornell-like potentials are employed as the central potentials, spin-spin and spin-orbit coupling derived from Breit-Fermi interaction are treated as hyperfine corrections, and model parameters are taken from previous works. The ground state $1^{-+}$ P-wave tetraquarks are predicted at 1.9, 4.2, and 6.6 GeV for the light, charmonium-like, and fully-charm sectors, respectively.
The decay width ratios of $1^{-+}$ tetraquark states are calculated for possible strong two-body decay channels within the rearrangement mechanism, including $\omega h_{1}$ and $\eta f_{1}$ for isospin $I=0$ light tetraquarks; $\rho h_{1}$ and $\pi f_{1}$ for isospin $I=1$ light tetraquarks; $\pi/\eta+\chi_{c1}$ and $\rho/\omega+h_{c}$ for charmonium-like tetraquarks; and $\eta_{c}\chi_{c1}$ and $J/\psi h_{c}$ for fully-charm tetraquarks. The theoretical results are compared with the observed exotic $1^{-+}$ states, and possible channels for observing $1^{-+}$ tetraquark states are also discussed. The work suggests that $\eta_{1}(1855)$ is unlikely to be a compact tetraquark state.
@@ I Introduction
Charmonium-like states with total angular momentum $J=1$ have been widely searched over the last two decades and are promising P-wave tetraquark candidates. Several exotic states in the charmonium region, known as Y states with $J^{PC}=1^{--}$ quantum numbers that mix with conventional $\psi$ mesons, could be P-wave tetraquark candidates: Y(4230) Aubert et al. (2005), Y(4360) Wang et al. (2007), Y(4660) Lees et al. (2014), the more recent Y(4484) Ablikim et al. (2022a), and Y(4544) Ablikim et al. (2024).
In conventional mesons composed of a quark and an antiquark $q\bar{q}$, the quantum numbers $J^{PC}$ follow simple rules: parity $P=(-1)^{L+1}$ and charge conjugation $C=(-1)^{L+S}$, where L is orbital angular momentum and S is total spin Navas et al. (2024). These rules allow common states such as $1^{--}$ (the $\rho$ meson) or $2^{++}$ (the $f_{2}$ meson) but not $1^{-+}$. Thus, a $1^{-+}$ state, once observed, is naturally considered to have a nonconventional structure, such as a hybrid meson, hadronic molecule, or tetraquark.
@@
The VES (Vertex Spectrometer) experiment observed the exotic $\pi_{1}(1400)$ state ($J^{PC}=1^{-+}$) in $\pi^{-}p\to\pi^{0}\eta n$ decays at 100 GeV Alde et al. (1988), with a mass of $1406\pm 20$ MeV and a width of $180\pm 30$ MeV. The analysis of the $\eta\pi^{0}$ channel showed a clear exotic signal inconsistent with conventional $q\bar{q}$ states.
The $\pi(1600)$ state was first observed by the Brookhaven E852 experiment Adams et al. (1998), which identified an exotic $1^{-+}$ resonance in the $\rho\pi$ channel. The analysis found a mass of $1593\pm 8^{+29}{-47}$ MeV and a width of $168\pm 20^{+150}{-12}$ MeV. Later, E852 Kuhn et al. (2004) observed two exotic $J^{PC}=1^{-+}$ resonances in $\pi^{-}p\to f_{1}(1285)\pi^{-}p$ collisions at 18 GeV: the known $\pi_{1}(1600)$ with mass $M=1709\pm 24\pm 41$ MeV and width $\Gamma=403\pm 80\pm 115$ MeV, and a new $\pi_{1}(2015)$ state at $M=2001\pm 30\pm 92$ MeV with width $\Gamma=333\pm 52\pm 49$ MeV.
In the updated PDG review, $\pi(1400)$ and $\pi(1600)$ are suggested to be a single state, as coupled-channel analyses favor one broad isovector resonance in the 1.4–1.6 GeV region Navas et al. (2024). Of the reported candidates, the $\pi_{1}(2015)$ has only been observed by the E852 experiment Kuhn et al. (2004), and its existence remains inconclusive.
The BESIII Collaboration recently observed the isoscalar exotic state $\eta_{1}(1855)$ with $J^{PC}=1^{-+}$ in $J/\psi\to\gamma\eta\eta^{\prime}$ decays Ablikim et al. (2022b), measuring its mass as $(1855\pm 9^{+6}_{-1})$ MeV, its width as $(188\pm 18^{+3}_{-8})$ MeV with $19\sigma$ significance, and the product branching fraction $\mathcal{B}(J/\psi\to\gamma\eta_{1}(1855))\times\mathcal{B}(\eta_{1}(1855)\to\eta\eta^{\prime})=(2.70\pm 0.41^{+0.16}_{-0.35})\times 10^{-6}$. The narrow width of $\eta_{1}(1855)$ contrasts sharply with the broad $\pi_{1}(1600)$, suggesting different underlying dynamics between these exotic states.
@@
The paper is organized as follows. In Sec. II, all possible configurations of the color, spin, flavor, and spatial degrees of freedom of tetraquark states are introduced, and a constituent quark model developed in our previous work Zhao et al. (2021, 2022, 2025) is briefly reviewed. In Sec. III, the $1^{-+}$ P-wave tetraquark mass spectra and decay branching ratios are evaluated within the constituent quark model. The theoretical results are compared with experimental data, and discussions of $\pi_{1}(2015)$ and $\eta_{1}(1855)$ are provided. A summary is given in Sec. IV.
@@ II.1 Quark configurations of tetraquark
In this work, we consider $J^{PC}=1^{-+}$ light ($q_{1}q_{2}\bar{q}_{3}\bar{q}_{4}$), charmonium-like ($q_{1}c_{2}\bar{q}_{3}\bar{c}_{4}$), and fully-charm ($c_{1}c_{2}\bar{c}_{3}\bar{c}_{4}$) tetraquark states, where $q$ and $c$ denote a light quark and a charm quark, and $\bar{q}$ and $\bar{c}$ denote a light antiquark and a charm antiquark, respectively.
The permutation symmetry of the two-quark cluster ($qq$, qcqc, or $cc$) in tetraquark states is described by the Young tableaux $2]_{6}$ and $11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{q}$, $\bar{q}\bar{c}$, or $\bar{c}\bar{c}$) consists of a $211]_{3}$ triplet and a $22]_{\bar{6}}$ antisextet.
@@
The $J^{PC}=1^{-+}$ tetraquark states correspond to $L=1$, $S=1$ configurations according to the parity ($P=(-1)^{L}$) and charge-conjugation ($C=(-1)^{L+S}$) quantum numbers, where $L$ is the total orbital angular momentum and $S$ is the total spin of the tetraquark. Thus, the $J^{PC}=1^{-+}$ tetraquark requires an odd $L$, with the lowest possible value being $L=1$. The complete basis is constructed by coupling harmonic-oscillator wave functions, which takes the general form:
@@
$u_{3}=\frac{(m_{1}+m_{2})+(m_{3}+m_{4})}{m_{1}+m_{2}+m_{3}+m_{4}},$ (3)
The complete bases of the tetraquarks with $N=1$ are listed in Table 1 for reference. The principal quantum number can be increased for each basis to expand the basis size. In this work, basis states with $N\leq 13$ are used to calculate the mass spectrum.
Table 1: The complete bases of tetraquark with quantum number, $N=2n+L$ and $L=1$.
@@
For constructing $S=1$ configuration, the possible spin combinations are $\left[\psi_{[s=1]}^{q_{1}q_{2}}\otimes\psi_{[s=1]}^{\bar{q}_{3}\bar{q}_{4}}\right]_{S=1}$, $\psi_{[s=1]}^{q_{1}q_{2}}\otimes\psi_{[s=0]}^{\bar{q}_{3}\bar{q}_{4}}$, and $\psi_{[s=0]}^{q_{1}q_{2}}\otimes\psi_{[s=1]}^{\bar{q}_{3}\bar{q}_{4}}$.
Because a $qq\bar{q}\bar{q}$ or $cc\bar{c}\bar{c}$ tetraquark must be a color singlet and simultaneously antisymmetric under any permutation of identical quarks, the $qq$-cluster spatial-spin-flavor configurations for $qq\bar{q}\bar{q}$ states are $\psi^{osf}_{[11]}$ for the color configuration $\psi^{c}_{[2]}$ and $\psi^{osf}_{[2]}$ for $\psi^{c}_{[11]}$. All possible configurations are as follows:
@@
Table 2: $I=0$ and $I=1$ Light tetraquark configurations for $J^{PC}=1^{-+}$ in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$. Pairwise Pauli is enforced for $(12)$ and $(34)$.
@@
Table 3: Possible fully-charm tetraquark configurations for $J^{PC}=1^{-+}$ in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$.
@@
For $qc\bar{q}\bar{c}$ tetraquarks, there is no Pauli-antisymmetry constraint tying flavor to color, spin, or space, so no pair contains two identical quarks or two identical antiquarks. The color-spin-orbital (CSO) sector can therefore mix freely, independent of flavor, and charmonium-like tetraquark states are isospin-blind. In this work, we use the notation $\psi^{c}_{\bar{3}\otimes 3}\psi^{s=1}_{(1\otimes 1)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{s=1}_{(1\otimes 0)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{s=1}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{s=1}_{(1\otimes 0)}$ to denote these configurations. The possible configurations of charmonium-like tetraquark are listed in Table 4
Table 4: Charmonium-like tetraquark configurations for $J^{PC}=1^{-+}$ in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$.
@@ II.2 Hamiltonian
We employ a widely used nonrelativistic Hamiltonian Wang et al. (2021); Wu and Zhu (2025); Zhao et al. (2025) to study meson and tetraquark systems, which takes the following form:
@@
where $M^{ave}_{k}$ denotes the spin-averaged mass as $\frac{1}{4}M_{PS}+\frac{3}{4}M_{V}$. $m_{k}$ are the constituent quark masses. $\vec{\lambda}_{i}$ are the quark color operator in SU(3). The central potential $V_{0}(r_{ij})$, the spin-spin interaction $V_{ss}(r_{ij})$ and spin-orbital interaction $V_{so}(r_{ij})$ are taking the form:
@@
$\quad-\frac{1}{r_{ij}}\frac{dV_{S}}{d(r_{ij})}\left(\frac{\vec{L_{ij}}\cdot\vec{s_{i}}}{2m_{i}^{2}}+\frac{\vec{L_{ij}}\cdot\vec{s_{i}}}{m_{j}^{2}}\right),$ (9)
@@
In line with the previous works Zhao et al. (2021, 2022, 2025), $A_{ij}$, $B_{ij}$, and $\sigma_{ij}$ are proposed to be mass dependent coupling parameters, taking the form
@@
The coupling constants $a$, $b$, $B_{0}$, and $\sigma_{0}$, as well as the constituent quark masses $m_{u}$ and $m_{c}$, are taken from previous work Zhao et al. (2021, 2025). These parameters were fixed by reproducing the S- and P-wave mass spectra of light, charmed, bottom, charmonium, and bottomonium mesons within the Hamiltonian in Eq. (II.2). The imported parameters and constituent quark masses are listed in Table 5, and the corresponding fits can be found in Ref. Zhao et al. (2025).
@@ II.3 Two-body strong decay
The strong two-body decay properties of the bound states, specifically the Y tetraquark states within the rearrangement mechanism, are studied, as shown in Figure 1.
@@
where $\psi_{Y}^{c}$, $\psi_{Y}^{s}$, and $\psi_{Y}^{f}$ are color, spin, and flavor wave functions of the initial states corresponding to the Y tetraquark states in mass spectrum. $\psi^{c}_{M_{1}M_{2}}$, $\psi^{s}_{M_{1}M_{2}}$ and $\psi^{f}_{M_{1}M_{2}}$ are the color, spin, and flavor wave functions of the final states of all possible two-body strong decay channels: $\omega h_{1}$ and $\eta f_{1}$ for isospin $I=0$ light tetraquarks; $\rho h_{1}$ and $\pi f_{1}$ for isospin $I=1$ light tetraquarks; $\pi/\eta+\chi_{c1}$ and $\rho/\omega+h_{c}$ for charmonium-like tetraquarks; and $\eta_{c}\chi_{c1}$ and $J/\psi h_{c}$ for fully-charm tetraquarks.
In the work, the ratio between $|T_{csf}|^{2}$ is considered to be branching ratio approximately supposing that the spatial part from different decay channels provide similar contribution.
Figure 1: Quark rearranged diagram for $1^{-+}$ tetraquark.

### v2

# Systematic study of exotic $1^{-+}$ tetraquark spectroscopy

@@ Abstract
The masses of exotic quantum-number $1^{-+}$ compact tetraquark states are calculated in a constituent quark model, where a Cornell-like potential is employed as the central potential, spin-spin and spin-orbit coupling derived from the Breit-Fermi interaction are treated as hyperfine corrections, and model parameters are taken from previous works. The ground state $1^{-+}$ P-wave tetraquarks are predicted at 1.9, 4.2, and 6.6 GeV for the light, charmonium-like, and fully-charm sectors, respectively. The decay width ratios of $1^{-+}$ tetraquark states are calculated for two-body strong decay channels within the rearrangement mechanism, including $\omega h_{1}$ and $\eta f_{1}$ for isospin $I=0$ light tetraquarks, $\rho h_{1}$ and $\pi f_{1}$ for isospin $I=1$ light tetraquarks, $\pi/\eta+\chi_{c1}$ and $\rho/\omega+h_{c}$ for charmonium-like tetraquarks, and $\eta_{c}\chi_{c1}$ and $J/\psi h_{c}$ for fully-charm tetraquarks. The theoretical results are compared with the observed exotic $1^{-+}$ states, and promising search channels for $1^{-+}$ tetraquarks are discussed. The work suggests that $\eta_{1}(1855)$ is unlikely to be a compact tetraquark state.
@@ I Introduction
Charmonium-like states with total angular momentum $J=1$ have been widely searched over the last two decades and are promising P-wave tetraquark candidates. Several exotic states in the charmonium region, known as Y states with $1^{--}$ quantum numbers that mix with conventional $\psi$ mesons, could be P-wave tetraquark candidates: Y(4230) Aubert et al. (2005), Y(4360) Wang et al. (2007), Y(4660) Lees et al. (2014), the more recent Y(4484) Ablikim et al. (2022a), and Y(4544) Ablikim et al. (2024).
In conventional mesons composed of a quark and an antiquark $q\bar{q}$, the quantum numbers $J^{PC}$ follow simple rules: parity $P=(-1)^{L+1}$ and charge conjugation $C=(-1)^{L+S}$, where L is orbital angular momentum and S is total spin Navas et al. (2024). These rules allow common states such as $1^{--}$ (the $\rho$ meson), $1^{+-}$ (the $h_{1}$ meson), or $1^{++}$ (the $f_{1}$ meson) but not $1^{-+}$. Thus, a $1^{-+}$ state, once observed, is naturally considered to have a nonconventional structure, such as a hybrid meson, hadronic molecule, or tetraquark.
@@
The VES (Vertex Spectrometer) experiment observed the exotic $1^{-+}$ $\pi_{1}(1400)$ state in $\pi^{-}p\to\pi\eta n$ decays at 100 GeV Alde et al. (1988), with a mass of $1406\pm 20$ MeV and a width of $180\pm 30$ MeV. The analysis of the $\eta\pi$ channel showed a clear exotic signal inconsistent with conventional $q\bar{q}$ states.
The $\pi(1600)$ state was first observed by the Brookhaven E852 experiment Adams et al. (1998), which identified an exotic $1^{-+}$ resonance in the $\rho\pi$ channel. The analysis found a mass of $1593\pm 8^{+29}_{-47}$ MeV and a width of $168\pm 20^{+150}_{-12}$ MeV. Later, E852 Kuhn et al. (2004) observed two exotic $1^{-+}$ resonances in $\pi^{-}p\to f_{1}(1285)\pi^{-}p$ collisions at 18 GeV: the known $\pi_{1}(1600)$ with mass $M=1709\pm 24\pm 41$ MeV and width $\Gamma=403\pm 80\pm 115$ MeV, and a new $\pi_{1}(2015)$ state at $M=2001\pm 30\pm 92$ MeV with width $\Gamma=333\pm 52\pm 49$ MeV.
In the updated Particle Data Group (PDG) review, $\pi(1400)$ and $\pi(1600)$ are suggested to be a single state, as coupled-channel analyses favor one broad isovector resonance in the 1.4–1.6 GeV region Navas et al. (2024). Of the reported candidates, the $\pi_{1}(2015)$ has only been observed by the E852 experiment Kuhn et al. (2004), and its existence remains inconclusive.
The BESIII Collaboration recently observed the isoscalar exotic state $\eta_{1}(1855)$ with $1^{-+}$ in $J/\psi\to\gamma\eta\eta^{\prime}$ decays Ablikim et al. (2022b), measuring its mass as $(1855\pm 9^{+6}_{-1})$ MeV, its width as $(188\pm 18^{+3}_{-8})$ MeV with $19\sigma$ significance, and the product branching fraction $\mathcal{B}(J/\psi\to\gamma\eta_{1}(1855))\times\mathcal{B}(\eta_{1}(1855)\to\eta\eta^{\prime})=(2.70\pm 0.41^{+0.16}_{-0.35})\times 10^{-6}$. The narrow width of $\eta_{1}(1855)$ contrasts sharply with the broad $\pi_{1}(1600)$, suggesting different underlying dynamics between these exotic states.
@@
The paper is organized as follows. In Sec. II, all possible configurations of the color, spin, flavor, and spatial degrees of freedom of tetraquark states are introduced, and a constituent quark model developed in our previous work Zhao et al. (2021, 2022, 2025) is briefly reviewed. In Sec. III, the $1^{-+}$ P-wave tetraquark mass spectra and relative decay ratios are evaluated within the constituent quark model. The theoretical results are compared with experimental data, and discussions of $\pi_{1}(2015)$ and $\eta_{1}(1855)$ are provided. A summary is given in Sec. IV.
@@ II.1 Quark configurations of tetraquark
In this work, we consider $1^{-+}$ light ($q_{1}q_{2}\bar{q}_{3}\bar{q}_{4}$), charmonium-like ($q_{1}c_{2}\bar{q}_{3}\bar{c}_{4}$), and fully-charm ($c_{1}c_{2}\bar{c}_{3}\bar{c}_{4}$) tetraquark states, where $q$ and $c$ denote a light quark and a charm quark, and $\bar{q}$ and $\bar{c}$ denote a light antiquark and a charm antiquark, respectively.
We employ a diquark-antidiquark coupling scheme and classify each cluster by SU(3) color Young tableaux. The permutation symmetry of the two-quark cluster ($qq$, qcqc, or $cc$) in tetraquark states is described by the Young tableaux $2]_{6}$ and $11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{q}$, $\bar{q}\bar{c}$, or $\bar{c}\bar{c}$) consists of a $211]_{3}$ triplet and a $22]_{\bar{6}}$ antisextet.
@@
The $1^{-+}$ tetraquark states correspond to $L=1$, $S=1$ configurations according to the parity ($P=(-1)^{L}$) and charge-conjugation ($C=(-1)^{L+S}$) quantum numbers, where $L$ is the total orbital angular momentum and $S$ is the total spin of the tetraquark. Thus, the $1^{-+}$ tetraquark requires an odd $L$, with the lowest possible value being $L=1$. The complete basis is constructed by coupling harmonic-oscillator wave functions, which takes the general form:
@@
$u_{3}=\frac{(m_{1}+m_{2})(m_{3}+m_{4})}{m_{1}+m_{2}+m_{3}+m_{4}},$ (3)
The size of complete bases can be expanded by increasing the principal quantum number $N$. The complete bases with $N=1$ are listed in Table 1 for reference. In this work, we employ complete bases of $N\leq 13$ to calculate the mass spectrum.
Table 1: The complete bases of tetraquark with the lowest quantum number, $N=2n+L=1$ and $L=1$.
@@
For $S=1$ configurations, the possible spin combinations are $\left[\psi_{[s=1]}^{q_{1}q_{2}}\otimes\psi_{[s=1]}^{\bar{q}_{3}\bar{q}_{4}}\right]_{S=1}$, $\psi_{[s=1]}^{q_{1}q_{2}}\otimes\psi_{[s=0]}^{\bar{q}_{3}\bar{q}_{4}}$, and $\psi_{[s=0]}^{q_{1}q_{2}}\otimes\psi_{[s=1]}^{\bar{q}_{3}\bar{q}_{4}}$. As a $qq\bar{q}\bar{q}$ or $cc\bar{c}\bar{c}$ tetraquark must be a color singlet and simultaneously antisymmetric under any permutation of identical quarks, the $qq$-cluster spatial-spin-flavor configurations for $qq\bar{q}\bar{q}$ states are $\psi^{osf}_{[11]}$ for the color configuration $\psi^{c}_{[2]}$ and $\psi^{osf}_{[2]}$ for $\psi^{c}_{[11]}$. All possible configurations are as follows:
@@
Table 2: $I=0$ and $I=1$ $1^{-+}$ Light tetraquark configurations in a complete harmonic oscillator (HO) basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$. Pairwise Pauli is enforced for $(12)$ and $(34)$.
@@
Table 3: Possible $1^{-+}$ fully-charm tetraquark configurations for $1^{-+}$ in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$.
@@
For $qc\bar{q}\bar{c}$ tetraquarks, there is no Pauli-antisymmetry constraint tying flavor to color, spin, or space, so no pair contains two identical quarks or two identical antiquarks. The color-spin-orbit (CSO) sector can therefore mix freely, independent of flavor, and charmonium-like tetraquark states are isospin-blind. In this work, we use the notation $\psi^{c}_{\bar{3}\otimes 3}\psi^{s=1}_{(1\otimes 1)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{s=1}_{(1\otimes 0)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{s=1}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{s=1}_{(1\otimes 0)}$ to denote these configurations. The possible configurations of charmonium-like tetraquark are listed in Table 4
Table 4: Possible $1^{-+}$ charmonium-like tetraquark configurations in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$.
@@ II.2 Hamiltonian
We employ a widely used nonrelativistic Hamiltonian Wang et al. (2021); Wu and Zhu (2025); Zhao et al. (2025) to study meson and tetraquark systems, which takes the form:
@@
where $M^{ave}_{k}$ denotes the spin-averaged mass as $\frac{1}{4}M_{PS}+\frac{3}{4}M_{V}$. $m_{k}$ are the constituent quark masses. $\vec{\lambda}_{i}$ are the quark color operator in SU(3). The central potential $V_{0}(r_{ij})$, the spin-spin interaction $V_{ss}(r_{ij})$ and spin-orbit interaction $V_{so}(r_{ij})$ take the form:
@@
$\quad-\frac{1}{r_{ij}}\frac{dV_{S}}{d(r_{ij})}\left(\frac{\vec{L_{ij}}\cdot\vec{s_{i}}}{2m_{i}^{2}}+\frac{\vec{L_{ij}}\cdot\vec{s_{j}}}{2m_{j}^{2}}\right),$ (9)
@@
In line with the previous works Zhao et al. (2021, 2022, 2025), $A_{ij}$, $B_{ij}$, and $\sigma_{ij}$ are taken to be mass-dependent coupling parameters, taking the form
@@
The coupling constants $a$, $b$, $B_{0}$, and $\sigma_{0}$, as well as the constituent quark masses $m_{u}$ and $m_{c}$, are taken from previous work Zhao et al. (2021, 2025). These parameters were fixed by reproducing the S- and P-wave mass spectra of light, charmed, bottom, charmonium, and bottomonium mesons within the Hamiltonian in Eq. (II.2). The adopted parameters and constituent quark masses are listed in Table 5, and the corresponding fits can be found in Ref. Zhao et al. (2025).
@@ II.3 Two-body strong decay
Figure 1: Quark rearranged diagram for $1^{-+}$ tetraquark.
We study the two-body strong decay properties of $1^{-+}$ Y tetraquark states within the rearrangement mechanism, as shown in Figure 1, which is believed to be the dominant mechanism for tetraquarks decaying into two mesons. The transition matrix of the color-spin-flavor takes the forms,
@@
where $\psi_{Y}^{c}$, $\psi_{Y}^{s}$, and $\psi_{Y}^{f}$ are color, spin, and flavor wave functions of the initial tetraquark states taken from the mass spectrum calculation. $\psi^{c}_{M_{1}M_{2}}$, $\psi^{s}_{M_{1}M_{2}}$ and $\psi^{f}_{M_{1}M_{2}}$ are the color, spin, and flavor wave functions of the final states of possible two-body strong decay channels: $\omega h_{1}$ and $\eta f_{1}$ for isospin $I=0$ light tetraquarks; $\rho h_{1}$ and $\pi f_{1}$ for isospin $I=1$ light tetraquarks; $\pi/\eta+\chi_{c1}$ and $\rho/\omega+h_{c}$ for charmonium-like tetraquarks; and $\eta_{c}\chi_{c1}$ and $J/\psi h_{c}$ for fully-charm tetraquarks.
To partially account for kinematic effects beyond the color–spin–flavor overlap, we further weight each channel by the corresponding two-body phase space evaluated at the tetraquark mass, and estimate the relative partial widths as
$\Gamma_{\alpha}\propto\Phi_{\alpha}(M_{Y})\,|T_{csf}^{(\alpha)}|^{2},$
$\Phi_{\alpha}(M_{Y})=k_{\alpha}\,\frac{E_{1}(k_{\alpha})\,E_{2}(k_{\alpha})}{M_{Y}},$ (12)
where $\alpha$ labels a given decay mode $Y\to M_{1}M_{2}$, $M_{Y}$ is the mass of the initial tetraquark, and $E_{1,2}(k)=\sqrt{m_{1,2}^{2}+k^{2}}$ are the on-shell energies of the final mesons with masses $m_{1,2}$. The breakup momentum $k_{\alpha}$ in the $Y$ rest frame is
$k_{\alpha}=\frac{\lambda^{1/2}(M_{Y}^{2},m_{1}^{2},m_{2}^{2})}{2M_{Y}},\qquad$ (13)
where $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$.
In the work, the ratios between $\Gamma_{\alpha}$ of various decay channels are approximated as the decay branching ratios (relative decay ratios), supposing that the contributions of the spatial part of different decay channels are not significantly different.